# Neural Networks for Extreme Quantile Regression with an Application to Forecasting of Flood Risk

### v3

@@ 3 Extreme Quantile Regression Neural Networks
In this section, we propose a new methodology that combines the extrapolation power of the GPD model with the high-dimensional predictor space capabilities and flexibility of neural networks to obtain accurate estimates for quantile functions $Q_{\mathbf{x}}(\tau)$ at extreme levels $\tau$. Let $\mathcal{D}=\left\{(\mathbf{x}_{i},y_{i})\right\}_{i=1}^{n}$ be the training dataset. Estimation of conditional extreme quantiles $\hat{Q}_{\mathbf{x}}(\tau)$ using (6) requires estimators for the intermediate quantile function $\hat{Q}_{\mathbf{x}}(\tau_{0})$ with $\tau_{0}<\tau$ and the GPD parameter $\sigma(\mathbf{x})$ and $\xi(\mathbf{x})$. It is customary to proceed in two steps.
@@
The intermediate probability $\tau_{0}$ should be chosen low enough to allow for stable estimation of $Q_{\mathbf{x}}(\tau_{0})$ with classical empirical methods, but high enough for the approximation in (5) to be accurate, so that the exceedances $z_{i}$ are approximate samples of a GPD. However, it is not a classical tuning parameter, since different values for $\tau_{0}$ yield different subsets of exceedances $\mathcal{I}$. Comparison of the loss function (11) on these datasets would therefore not be meaningful. Instead, the threshold is, in the univariate case, usually selected in terms of stability plots and sensitivity analyses.
In the second step, we estimate the GPD parameters $\sigma(\mathbf{x})$ and $\xi(\mathbf{x})$ based on the set of exceedances $z_{i}$, $i\in\mathcal{I}$. Modelling these parameters directly in the extrapolation formula (6) may lead to strong dependence between the estimates and numerical instabilities. We therefore rely on an orthogonal reparametrization that has a diagonal Fisher information matrix. As for the standard asymptotic GPD likelihood properties, the latter is well-defined for the GPD model when $\xi(\mathbf{x})>-0.5$, and the reparametrization
$(\sigma(\mathbf{x}),\xi(\mathbf{x}))\mapsto(\nu(\mathbf{x}),\xi(\mathbf{x})),\qquad\nu(\mathbf{x}):=\sigma(\mathbf{x})(\xi(\mathbf{x})+1),$
yields the desired orthogonality (Cox and Reid, 1987; Chavez-Demoulin and Davison, 2005). In our experiments, this reparametrization significantly improves stability and convergence in every considered setting.
In this section, we propose a flexible neural network model for the orthogonalized GPD parameters $\nu(\mathbf{x};\mathcal{W})$ and $\xi(\mathbf{x};\mathcal{W})$, where $\mathcal{W}$ denotes the collection of all model parameters. This can be seen as conditional density estimation with output dimension $q=2$ where the parametric family is the GPD model with parameters $\theta=(\nu,\xi)$ depending on the covariate $\mathbf{X}=\mathbf{x}$. In general, an estimate of the model parameters $\hat{\mathcal{W}}$ is thus obtained as a minimizer of the GPD deviance loss over the training exceedances
@@
This yields a flexible model for the conditional tail distribution of $Y\mid\mathbf{X}=\mathbf{x}$ with which not only $\hat{Q}_{\mathbf{x}}(\tau)$ can be regressed, but also conditional exceedance probabilities over a high threshold or conditional expected shortfalls, for example.
@@ 3.1 Independent Observations
We first consider the case where the training data $\mathcal{D}=\left\{(\mathbf{x}_{i},y_{i})\right\}_{i=1}^{n}$ is a set of independent, identically distributed observations of $(\mathbf{X},Y)$. The goal in this case is the estimation of the conditional quantile $Q_{\mathbf{x}}(\tau)$ for a predictor value $\mathbf{X}=\mathbf{x}$ at an extreme level $\tau>0$. For the first step of estimating the intermediate quantile function, in principle, any classical quantile regression method can be used. To avoid overfitting and obtain unbiased generalization error estimates from the training set $\mathcal{D}$, the predicted $\hat{Q}_{\mathbf{x}_{i}}(\tau_{0})$, $i=1,\dots,n$, should be constructed out of training sample. This is achievable in two ways. Using bagging methods such as generalized random forests (Athey et al., 2019) is a convenient choice since they allow for out-of-bag predictions where only a single fit on $\mathcal{D}$ is required. For other methods such as quantile regression neural networks (Cannon, 2011), out of training sample predictions can be obtained in a foldwise manner similar to cross-validation. In the sequel, we assume that the intermediate quantile, and thus the exceedances, are given.
@@
The only restrictions are on the output activation functions, since ${\nu}(\mathbf{x})$ should be strictly positive. We find the exponential function or the ${\rm SELU}$ activation (Klambauer et al., 2017) shifted above zero to be good choices. Regarding the output activation for ${\xi}$, no strict restrictions apply and the identity would be a natural choice. However, standard likelihood regularity properties are not satisfied for the GPD model when $\xi\leq-0.5$, which very rarely occurs in practice. We observe that smoothly restricting the shape estimates, for example between $-0.5$ and $0.7$ with the activation $x\mapsto 0.6\tanh{(x)}+0.1$, helps to improve training stability. This avoids aberrant $\xi$ estimates in the early stages of the training and still covers almost all practical cases. In many situations, it is reasonable to assume that only the scale $\nu(\mathbf{x})$ varies locally but $\xi(\mathbf{x})\equiv\xi$ is constant (e.g., Kinsvater et al., 2016). This can be achieved by restricting the network so that the shape output only depends on a bias term.
@@
The tuning parameters $\Theta$ for the conditional GPD density estimation network $g_{\mathcal{W}}$ and the intermediate quantile model $\hat{Q}_{\cdot}(\tau_{0})$ capable of out-of-sample prediction are pre-specified. The training data $\mathcal{D}=\left\{(\mathbf{x}_{i},y_{i})\right\}_{i=1}^{n}$ and test covariates $\mathbf{x}$ are observed. Let $\tau\in(\tau_{0},1)$ be the desired probability level.
@@
The conditional GPD estimation in the second step relies on the exceedances to ensure that only information from the tail is used for extrapolation. There may however be residual information in the moderately extreme observations that should not be discarded. We propose to use the intermediate quantiles $\hat{Q}_{\mathbf{x}_{i}}(\tau_{0})$, as an additional feature in the conditional density estimation. This feature engineering seems to consistently and significantly improve the accuracy of the final prediction $\hat{Q}_{\mathbf{x}}(\tau)$ on test data in our simulations. This idea is to some degree related to stacked learning (Breiman, 1996; Wolpert, 1992) and is achieved by considering the $(p+1)$-dimensional vector $(\mathbf{x}_{i},\hat{Q}_{\mathbf{x}_{i}}(\tau_{0}))$, $i=1,\dots,n,$ instead of $\mathbf{x}_{i}$ as predictors, for the network input. To avoid overfitting, it is again important that the intermediate quantile estimates are constructed out of training sample. This new feature also improves other extreme quantile regression methods, such as the GBEX model (Velthoen et al., 2023), as observed in our simulations in Section 4.
@@
Since the true quantile $Q_{\mathbf{x}}(\tau)$ is unknown in real-world data, we can not assess the performance of $\hat{Q}_{\mathbf{x}}(\tau)$ with metrics such as mean squared or absolute errors. As illustrated in Section 2.1, the quantile loss is also unreliable due to the data scarcity at extreme quantiles. We therefore choose the final validation loss based on the GPD deviance to compare different choices of hyperparameters, as it is the most reliable surrogate metric.
@@ 3.2 Sequential Dependence
In many applications, the observations are not independent but display sequential dependence, such as in time series. In this case, we denote the training data by $\mathcal{D}=\left\{(\mathbf{x}_{t},y_{t})\right\}_{t=1}^{T}$, which are observed sequentially from a time series $\left\{(\mathbf{X}_{t},Y_{t})\right\}_{t=1}^{T}$. The goal here is different from the case of independent observations. Indeed, we would like to predict as well as possible high quantiles of the response $Y_{u}$ at some time point $u$ one step in the future based on all past information $\tilde{\mathbf{X}}_{u}:=\left\{(\mathbf{X}_{t},Y_{t})\right\}_{t<u}$. Therefore, the target is
@@
where $\tilde{\mathbf{x}}_{u}:=\left\{(\mathbf{x}_{t},y_{t})\right\}_{t<u}$ are observations that are not necessarily part of the training set. In this section, we propose a recurrent neural network to solve this task. Several principles are the same as in the case of independent observations, such as the choice of output activation functions. We thus focus on the differences to the independent case.
While in principle it is still possible to use any classical quantile regression method to model the intermediate conditional quantiles at level $\tau_{0}$, we recommend using quantile regression neural networks (Cannon, 2011; Zhang et al., 2019) with a recurrent architecture. These models are specifically designed for sequential dependence and can easily adapt to varying sequence lengths of the input features $\tilde{\mathbf{X}}_{t}$. In our experiments, recurrent quantile regression neural networks consistently outperformed generalized random forests in the presence of sequential dependence. Although a varying sequence length of input features is possible, we restrict this length to a fixed horizon $s\ll T$, for computational efficiency. Thus, for any time point $t$ we define its past by $\tilde{\mathbf{x}}_{t}=\{(\mathbf{x}_{j},y_{j})\}_{j=t-s}^{t-1}$. For simplicity, we denote our augmented training set by $\tilde{\mathcal{D}}=\{(\tilde{\mathbf{x}}_{t},y_{t})\}_{t=s+1}^{T}$.
@@
The tuning parameters $\Theta$ for the recurrent conditional GPD density estimation network $\tilde{g}_{\mathcal{W}}$, the intermediate quantile model $\hat{Q}_{\cdot}(\tau_{0})$ capable of out of sample prediction and horizon $s$ are pre-specified. The training data $\tilde{\mathcal{D}}=\left\{(\tilde{\mathbf{x}}_{t},y_{t})\right\}_{t=s+1}^{T}$ and test covariates $\tilde{\mathbf{x}}_{u}$ are observed. Let $\tau\in(\tau_{0},1)$ be the desired probability level.
@@
Algorithm 2 summarizes our EQRN for sequential data. For the estimation of the conditional GPD parameters, the main difference compared to the independent model is the use of a recurrent architecture to capture the sequential nature of the data. If validation splitting is used during training, the split should preserve the sequential structure instead of being performed randomly. For the use of the intermediate quantile as a feature, two approaches seem relevant. The first one is to only use $\hat{Q}_{\tilde{\mathbf{x}}_{t}}(\tau_{0})$ as a separate additional input to $\tilde{\mathbf{x}}_{t}$ in the network. The second approach is to also use past intermediate information by considering $\{(\mathbf{x}_{j},y_{j},\hat{Q}_{\tilde{\mathbf{x}}_{j}}(\tau_{0}))\}_{j=t-s}^{t-1}$ instead of $\tilde{\mathbf{x}}_{t}$ as input features to model the GPD parameters $\nu(\tilde{\mathbf{x}}_{t})$ and $\xi(\tilde{\mathbf{x}}_{t})$. We prefer the second approach as it can pass more information to the tail model. More details on the function calls in the algorithm can be found in Supplementary Material S.2.
The training data consists of time points $\{1,\dots,T\}$, while the test data uses information from time points $\{u-s,\dots,u-1\}$ to predict at time $u$. These two intervals are typically disjoint when the model was fitted in the past and is applied for prediction in the present. The prediction model can of course be used to predict on the training data when $u\leq T$, but such predictions might be overly precise since $y_{u}$ was used in the training procedure.
@@ 4.1 Setup
In this section, we assess the accuracy of our EQRN model in predicting extreme conditional quantiles on simulated data and compare it to existing state-of-the-art methods. The aim is to study a simplified version of the application, which motivates the simulation setup and modelling choices. We thus focus on the case of sequentially dependent data; a simulation study for independent data can be found in Supplementary Material S.3.
The main competitors from the extreme value literature in terms of flexibility are the generalized additive models (EGAM) (Youngman, 2019) and gradient boosting for extreme quantile regression (GBEX) (Velthoen et al., 2023), which both use conditional GPD modelling. For sequentially dependent data, we also consider the extreme quantile autoregression (EXQAR) (Li and Wang, 2019) as, although assuming linear quantile dependence, the model is designed for time series. As a benchmark, we consider an unconditional GPD model that ignores the covariate dependence altogether, and a semi-conditional GPD model that uses the covariate only for the intermediate quantile. We include results for the generalized random forests for quantile regression (GRF) (Athey et al., 2019), which does not use extrapolation for high quantiles. The training data $\mathcal{D}=\left\{(\mathbf{x}_{t},y_{t})\right\}_{t=1}^{T}$, with $T=\numprint{7000}$, are sequentially generated from the time series
@@
Figure S.5 in Supplementary Material S.4 shows part of the simulated data. To have a fair comparison, all methods use the same covariate vectors $\tilde{\mathbf{x}}_{t}=\{(\mathbf{x}_{j},y_{j})\}_{j=t-s}^{t-1}$ with $s=10$. This model admits a GPD approximation as in (5) since the conditional distribution $Y_{t}\mid\tilde{\mathbf{X}}_{t}=\tilde{\mathbf{x}}_{t}$ is a folded normal distribution, and its tail can therefore be approximated by a GPD with shape parameter $\xi(\tilde{\mathbf{x}}_{t})=0$. For the methods that use covariate-dependent intermediate quantiles, we use the same estimates $\hat{Q}_{\tilde{\mathbf{x}}_{t}}(\tau_{0})$ with $\tau_{0}=80\%$ from a recurrent quantile regression neural network (QRN); for a sensitivity analysis of the choice of $\tau_{0}$ see Supplementary Material S.4. The best QRN architecture and hyperparameters are chosen based on validation quantile loss. For the methods that use covariate-dependent GPD parameters, we also use $\hat{Q}_{\tilde{\mathbf{x}}_{t}}(\tau_{0})$ as additional covariate. Although designed for univariate time series, we adapted EXQAR to accept several covariate sequences. Those two choices significantly improve the competitors’ performances.

### v4

@@ 3 Extreme quantile regression neural networks
In this section we propose a new methodology that combines the extrapolation power of the GPD model with the high-dimensional predictor space capabilities and flexibility of neural networks to obtain accurate estimates for quantile functions $Q_{\mathbf{x}}(\tau)$ at extreme levels $\tau$. Let $\mathcal{D}=\left\{(\mathbf{x}_{i},y_{i})\right\}_{i=1}^{n}$ be the training dataset. Estimation of conditional extreme quantiles $\hat{Q}_{\mathbf{x}}(\tau)$, using (6), requires estimators for the intermediate quantile function $\hat{Q}_{\mathbf{x}}(\tau_{0})$ with $\tau_{0}<\tau$ and the GPD parameter $\sigma(\mathbf{x})$ and $\xi(\mathbf{x})$. It is customary to proceed in two steps.
@@
The intermediate probability $\tau_{0}$ should be chosen low enough to allow for stable estimation of $Q_{\mathbf{x}}(\tau_{0})$ with classical empirical methods, but high enough for the approximation in (5) to be accurate, so that the exceedances $z_{i}$ are approximate samples of a GPD. However, it is not a classical tuning parameter, since different values for $\tau_{0}$ yield different subsets of exceedances $\mathcal{I}$. Comparison of the loss function (11) on these datasets would, therefore, not be meaningful. Instead, the threshold is, in the univariate case, usually selected in terms of stability plots and sensitivity analyses.
In the second step, we estimate the GPD parameters $\sigma(\mathbf{x})$ and $\xi(\mathbf{x})$ based on the set of exceedances $z_{i}$, $i\in\mathcal{I}$. Modelling these parameters directly in the extrapolation formula (6) may lead to strong dependence between the estimates and numerical instabilities. We, therefore, rely on an orthogonal reparametrization that has a diagonal Fisher information matrix. As for the standard asymptotic GPD likelihood properties, the latter is well-defined for the GPD model, when $\xi(\mathbf{x})>-0.5$, and the reparametrization
$(\sigma(\mathbf{x}),\xi(\mathbf{x}))\mapsto(\nu(\mathbf{x}),\xi(\mathbf{x})),\qquad\nu(\mathbf{x}):=\sigma(\mathbf{x})(\xi(\mathbf{x})+1)$
yields the desired orthogonality (Cox and Reid, 1987; Chavez-Demoulin and Davison, 2005). In our experiments this reparametrization significantly improves stability and convergence in every considered setting.
In this section we propose a flexible neural network model for the orthogonalized GPD parameters $\nu(\mathbf{x};\mathcal{W})$ and $\xi(\mathbf{x};\mathcal{W})$, where $\mathcal{W}$ denotes the collection of all model parameters. This can be seen as conditional density estimation with output dimension $q=2$, where the parametric family is the GPD model with parameters $\theta=(\nu,\xi)$ depending on the covariate $\mathbf{X}=\mathbf{x}$. In general, an estimate of the model parameters $\hat{\mathcal{W}}$ is thus obtained as a minimizer of the GPD deviance loss over the training exceedances
@@
This yields a flexible model for the conditional tail distribution of $Y\mid\mathbf{X}=\mathbf{x}$ with which not only $\hat{Q}_{\mathbf{x}}(\tau)$ can be regressed but also conditional exceedance probabilities over a high threshold or conditional expected shortfalls, for example.
@@ 3.1 Independent observations
We first consider the case where the training data $\mathcal{D}=\left\{(\mathbf{x}_{i},y_{i})\right\}_{i=1}^{n}$ is a set of independent, identically distributed observations of $(\mathbf{X},Y)$. The goal in this case is the estimation of the conditional quantile $Q_{\mathbf{x}}(\tau)$ for a predictor value $\mathbf{X}=\mathbf{x}$ at an extreme level $\tau>0$. For the first step of estimating the intermediate quantile function, in principle, any classical quantile regression method can be used. To avoid overfitting and obtain unbiased generalization error estimates from the training set $\mathcal{D}$, the predicted $\hat{Q}_{\mathbf{x}_{i}}(\tau_{0})$, $i=1,\dots,n$, should be constructed out of training sample. This is achievable in two ways. Using bagging methods, such as generalized random forests (Athey et al., 2019), is a convenient choice since they allow for out-of-bag predictions where only a single fit on $\mathcal{D}$ is required. For other methods such as quantile regression neural networks (Cannon, 2011), out of training sample predictions can be obtained in a foldwise manner similar to cross-validation. In the sequel we assume that the intermediate quantile, and thus the exceedances, are given.
@@
The only restrictions are on the output activation functions, since ${\nu}(\mathbf{x})$ should be strictly positive. We find the exponential function or the ${\rm SELU}$ activation (Klambauer et al., 2017) shifted above zero to be good choices. Regarding the output activation for ${\xi}$, no strict restrictions apply and the identity would be a natural choice. However, standard likelihood regularity properties are not satisfied for the GPD model when $\xi\leq-0.5$, which very rarely occurs in practice. We observe that smoothly restricting the shape estimates, for example, between $-0.5$ and $0.7$ with the activation $x\mapsto 0.6\tanh{(x)}+0.1$, helps to improve training stability. This avoids aberrant $\xi$ estimates in the early stages of the training and still covers almost all practical cases. In many situations it is reasonable to assume that only the scale $\nu(\mathbf{x})$ varies locally but $\xi(\mathbf{x})\equiv\xi$ is constant (e.g., Kinsvater et al., 2016). This can be achieved by restricting the network so that the shape output only depends on a bias term.
@@
The tuning parameters $\Theta$ for the conditional GPD density estimation network $g_{\mathcal{W}}$ and the intermediate quantile model $\hat{Q}_{\cdot}(\tau_{0})$ capable of out-of-sample prediction are prespecified. The training data $\mathcal{D}=\left\{(\mathbf{x}_{i},y_{i})\right\}_{i=1}^{n}$ and test covariates $\mathbf{x}$ are observed. Let $\tau\in(\tau_{0},1)$ be the desired probability level.
@@
The conditional GPD estimation in the second step relies on the exceedances to ensure that only information from the tail is used for extrapolation. There may, however, be residual information in the moderately extreme observations that should not be discarded. We propose to use the intermediate quantiles $\hat{Q}_{\mathbf{x}_{i}}(\tau_{0})$ as an additional feature in the conditional density estimation. This feature engineering seems to consistently and significantly improve the accuracy of the final prediction $\hat{Q}_{\mathbf{x}}(\tau)$ on test data in our simulations. This idea is, to some degree, related to stacked learning (Breiman, 1996; Wolpert, 1992) and is achieved by considering the $(p+1)$-dimensional vector $(\mathbf{x}_{i},\hat{Q}_{\mathbf{x}_{i}}(\tau_{0}))$, $i=1,\dots,n,$ instead of $\mathbf{x}_{i}$ as predictors, for the network input. To avoid overfitting, it is again important that the intermediate quantile estimates are constructed out of training sample. This new feature also improves other extreme quantile regression methods, such as the GBEX model (Velthoen et al., 2023), as observed in our simulations in Section 4.
@@
Since the true quantile $Q_{\mathbf{x}}(\tau)$ is unknown in real-world data, we can not assess the performance of $\hat{Q}_{\mathbf{x}}(\tau)$ with metrics such as mean squared or absolute errors. As illustrated in Section 2.1, the quantile loss is also unreliable due to the data scarcity at extreme quantiles. We, therefore, choose the final validation loss based on the GPD deviance to compare different choices of hyperparameters, as it is the most reliable surrogate metric.
@@ 3.2 Sequential dependence
In many applications the observations are not independent but display sequential dependence, such as in time series. In this case we denote the training data by $\mathcal{D}=\left\{(\mathbf{x}_{t},y_{t})\right\}_{t=1}^{T}$, which are observed sequentially from a time series $\left\{(\mathbf{X}_{t},Y_{t})\right\}_{t=1}^{T}$. The goal here is different from the case of independent observations. Indeed, we would like to predict as well as possible high quantiles of the response $Y_{u}$ at some time point $u$ one step in the future based on all past information $\tilde{\mathbf{X}}_{u}:=\left\{(\mathbf{X}_{t},Y_{t})\right\}_{t<u}$. Therefore, the target is
@@
where $\tilde{\mathbf{x}}_{u}:=\left\{(\mathbf{x}_{t},y_{t})\right\}_{t<u}$ are observations that are not necessarily part of the training set. In this section we propose a recurrent neural network to solve this task. Several principles are the same as in the case of independent observations, such as the choice of output activation functions. We thus focus on the differences to the independent case.
While in principle it is still possible to use any classical quantile regression method to model the intermediate conditional quantiles at level $\tau_{0}$, we recommend using quantile regression neural networks (Cannon, 2011; Zhang et al., 2019) with a recurrent architecture. These models are specifically designed for sequential dependence and can easily adapt to varying sequence lengths of the input features $\tilde{\mathbf{X}}_{t}$. In our experiments, recurrent quantile regression neural networks consistently outperformed generalized random forests in the presence of sequential dependence. Although a varying sequence length of input features is possible, we restrict this length to a fixed horizon $s\ll T$ for computational efficiency. Thus, for any time point $t$, we define its past by $\tilde{\mathbf{x}}_{t}=\{(\mathbf{x}_{j},y_{j})\}_{j=t-s}^{t-1}$. For simplicity, we denote our augmented training set by $\tilde{\mathcal{D}}=\{(\tilde{\mathbf{x}}_{t},y_{t})\}_{t=s+1}^{T}$.
@@
The tuning parameters $\Theta$ for the recurrent conditional GPD density estimation network $\tilde{g}_{\mathcal{W}}$, the intermediate quantile model $\hat{Q}_{\cdot}(\tau_{0})$ capable of out of sample prediction and horizon $s$ are prespecified. The training data $\tilde{\mathcal{D}}=\left\{(\tilde{\mathbf{x}}_{t},y_{t})\right\}_{t=s+1}^{T}$ and test covariates $\tilde{\mathbf{x}}_{u}$ are observed. Let $\tau\in(\tau_{0},1)$ be the desired probability level.
@@
Algorithm 2 summarizes our EQRN for sequential data. For the estimation of the conditional GPD parameters, the main difference compared to the independent model is the use of a recurrent architecture to capture the sequential nature of the data. If validation splitting is used during training, the split should preserve the sequential structure instead of being performed randomly. For the use of the intermediate quantile as a feature, two approaches seem relevant. The first one is to only use $\hat{Q}_{\tilde{\mathbf{x}}_{t}}(\tau_{0})$ as a separate additional input to $\tilde{\mathbf{x}}_{t}$ in the network. The second approach is to also use past intermediate information by considering $\{(\mathbf{x}_{j},y_{j},\hat{Q}_{\tilde{\mathbf{x}}_{j}}(\tau_{0}))\}_{j=t-s}^{t-1}$, instead of $\tilde{\mathbf{x}}_{t}$, as input features to model the GPD parameters $\nu(\tilde{\mathbf{x}}_{t})$ and $\xi(\tilde{\mathbf{x}}_{t})$. We prefer the second approach, as it can pass more information to the tail model. More details on the function calls in the algorithm can be found in Supplementary Material S.2.
The training data consists of time points $\{1,\dots,T\}$, while the test data uses information from time points $\{u-s,\dots,u-1\}$ to predict at time $u$. These two intervals are typically disjoint when the model was fitted in the past and is applied for prediction in the present. The prediction model can, of course, be used to predict on the training data when $u\leq T$, but such predictions might be overly precise since $y_{u}$ was used in the training procedure.
@@ 4.1 Setup
In this section we assess the accuracy of our EQRN model in predicting extreme conditional quantiles on simulated data and compare it to existing state-of-the-art methods. The aim is to study a simplified version of the application, which motivates the simulation setup and modelling choices. We thus focus on the case of sequentially dependent data; a simulation study for independent data can be found in Supplementary Material S.3.
The main competitors from the extreme value literature in terms of flexibility are the generalized additive models (EGAM) (Youngman, 2019) and gradient boosting for extreme quantile regression (GBEX) (Velthoen et al., 2023), which both use conditional GPD modelling. For sequentially dependent data, we also consider the extreme quantile autoregression (EXQAR) (Li and Wang, 2019) as, although assuming linear quantile dependence, the model is designed for time series. As a benchmark we consider an unconditional GPD model that ignores the covariate dependence altogether and a semiconditional GPD model that uses the covariate only for the intermediate quantile. We include results for the generalized random forests for quantile regression (GRF) (Athey et al., 2019), which does not use extrapolation for high quantiles. The training data $\mathcal{D}=\left\{(\mathbf{x}_{t},y_{t})\right\}_{t=1}^{T}$, with $T=\numprint{7000}$, are sequentially generated from the time series
@@
Figure S.5 in Supplementary Material S.4 shows part of the simulated data. To have a fair comparison, all methods use the same covariate vectors $\tilde{\mathbf{x}}_{t}=\{(\mathbf{x}_{j},y_{j})\}_{j=t-s}^{t-1}$ with $s=10$. This model admits a GPD approximation as in (5) since the conditional distribution $Y_{t}\mid\tilde{\mathbf{X}}_{t}=\tilde{\mathbf{x}}_{t}$ is a folded normal distribution, and its tail can, therefore, be approximated by a GPD with shape parameter $\xi(\tilde{\mathbf{x}}_{t})=0$. For the methods that use covariate-dependent intermediate quantiles, we use the same estimates $\hat{Q}_{\tilde{\mathbf{x}}_{t}}(\tau_{0})$ with $\tau_{0}=80\%$ from a recurrent quantile regression neural network (QRN); for a sensitivity analysis of the choice of $\tau_{0}$, see Supplementary Material S.4. The best QRN architecture and hyperparameters are chosen based on validation quantile loss. For the methods that use covariate-dependent GPD parameters, we also use $\hat{Q}_{\tilde{\mathbf{x}}_{t}}(\tau_{0})$ as additional covariate. Although designed for univariate time series, we adapted EXQAR to accept several covariate sequences. Those two choices significantly improve the competitors’ performances.